Find solution of equation by using the gaussian elimination method.

2x+3y+z=2

-x+2y+2z=1

x-y-2z=-1



1
Expert's answer
2021-02-04T08:05:21-0500

The augmented matrix for the set of equations is:

"\begin{bmatrix}
 2 & 3 & 1 & 2 \\
 -1 & 2 & 2 & 1 \\
 1 & -1 & -2 & -1
\end{bmatrix}"

Rearranging the rows:

"\begin{bmatrix}
 1 & -1 & -2 & -1 \\
 2 & 3 & 1 & 2 \\
 -1 & 2 & 2 & 1
\end{bmatrix}"

Multiply the first row by -2 and add it to the second row:

"\begin{bmatrix}
 1 & -1 & -2 & -1 \\
 0 & 5 & 5 & 4 \\
 -1 & 2 & 2 & 1
\end{bmatrix}"

Add the first row to the third row:

"\begin{bmatrix}
 1 & -1 & -2 & -1 \\
 0 & 5 & 5 & 4 \\
 0 & 1 & 0 & 0
\end{bmatrix}"

The third row translates to:

"y=0"

Therefore, the second row translates to:

"5z=4"

"z=4\/5"

The first row then translates to:

"x-2(4\/5)=-1"

"x-8\/5=-1"

"x=-1+8\/5"

"x=3\/5"